5 Find the maximum height and range for vo = 10 m/s, and ? = 90°.

5

Find the maximum height and range for vo = 10 m/s, and ? = 90°.

Grade:12

2 Answers

ruchi yadav
10 years ago
Maximum height in projectile motion ish={v_{0}^{2}\sin ^{2}(\theta ) \over {2g}}

h = (10)2sin2(900)/ 2(10)

h = 5 m


Range in projectile motion is d = \frac{v^2 \sin 2 \theta}{g}


d = (10)2 sin(1800)/10

Range =0
ruchi yadav
10 years ago
Maximum height in projectile motion ish={v_{0}^{2}\sin ^{2}(\theta ) \over {2g}}

h = (10)2sin2(900)/ 2(10)

h = 5 m


Range in projectile motion is d = \frac{v^2 \sin 2 \theta}{g}


d = (10)2sin(1800)/10

Range =0